# Hölder regularity of harmonic functions on metric measure spaces

### v2

@@ 2 Statement of the main results
Let us compare the stability of “gradient” and “Hölder” conditions as follows.
By a standard technique, it is easy to see that under $(UE)$, any one of the equivalent “gradient” conditions implies the heat kernel near-diagonal lower Gaussian bound $(NLE)$. In particular, under $(UE)$, we have ($\text{GLY}_{\infty}$) implies $(NLE)$. However, the converse is usually not true. Indeed, the conjunction of $(UE)$ and $(NLE)$, or equivalently, HK(2), is equivalent to the conjunction of the volume doubling condition and the scale-invariant $L^{2}$-Poincaré inequality on balls, hence is quasi-isometry stable. However, to obtain ($\text{GLY}_{\infty}$), some curvature assumptions are usually needed, see [47, 16], hence ($\text{GLY}_{\infty}$) is quasi-isometry unstable.
For comparison, by our main result, Theorem 2.1, under UHK($\Psi$), all the “Hölder” conditions are equivalent to NLE($\Psi$). By the quasi-isometry stability of UHK($\Psi$) (see [2]) and the conjunction of UHK($\Psi$) and NLE($\Psi$) (see [31]), our “Hölder” conditions are quasi-isometry stable. Roughly speaking, the stability originates from the strongly recurrent condition, which requires our MMD spaces to behave like $\mathbb{R}^{1}$ in the classical Riemannian setting.
@@
$(X,d,m,\mathcal{E},\mathcal{F})$ be the corresponding unbounded cable system with an elementary gradient operator $\nabla$, see [25, Section 3] for more details. Let $\alpha_{1}=1$, $\beta_{1}=2$, $\alpha_{2}=\alpha$, $\beta_{2}=\beta$.
@@
$\lVert\lvert\nabla u\rvert\rVert_{L^{\infty}(B)}\leq C_{H}\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$ GRH($\Phi,\Psi$)
It is obvious that on an unbounded cable system, HR($\Phi,\Psi$) implies GRH($\Phi,\Psi$). Indeed, assume HR($\Phi,\Psi$), then the LHS of HR($\Phi,\Psi$) becomes $|u(x)-u(y)|/d(x,y)$ for $d(x,y)\in(0,1)$. Letting $y\to x$ and using the Lebesgue density theorem, we obtain the limit $|\nabla u(x)|$. Taking essential supremum with respect to $x\in B$, we have GRH($\Phi,\Psi$). Roughly speaking, HR($\Phi,\Psi$) encodes more Hölder continuity information than GRH($\Phi,\Psi$) about the behavior of harmonic functions at large scale.
We give some results as important applications of our main result, Theorem 2.1, on unbounded cable systems. Firstly, due to the fact that two-sided heat kernel estimates have already obtained on a large class of fractals (see [5, 34, 42, 43]), including the Sierpiński carpet (see [8, 9, 35]), we have GRH($\Phi,\Psi$) holds on a large class of fractal-like cable systems, including the Sierpiński carpet cable system, whose validity was left open in [25].
@@
HK($\Psi$) with $\alpha<\beta$. Then GRH($\Phi,\Psi$) holds. In particular, GRH($\Phi,\Psi$) holds on the Sierpiński carpet cable system with $\alpha=\log 8/\log 3<2<\beta\approx 2.09697$.
@@
In [25], the authors verified directly GRH($\Phi,\Psi$) on the Vicsek cable system and the Sierpiński gasket cable system using certain Hölder regularity  [^1] The Hölder regularity follows indeed from certain extension algorithm for harmonic functions. of harmonic functions on the corresponding fractals [52], which is also available on some class of p.c.f. self-similar sets [54] without any technical difficulty, but not available on general fractals, especially on the Sierpiński carpet.
Secondly, we improve the results about gradient estimate for heat kernel as follows.
@@
HK($\Psi$) with $\alpha<\beta$.
@@
(1) is certain improvement of [25, Theorem 1.1] and (2) is certain improvement of [25, Corollary 1.3], where GRH($\Phi,\Psi$) was needed. Now we can obtain gradient estimate result GHK($\Phi,\Psi$) without assuming gradient type conditions like GRH($\Phi,\Psi$).
As a direct consequence of $\text{HHK}_{\text{exp}}(\Phi,\Psi)$, we also obtain the following Hölder estimate for the heat kernel on unbounded MMD spaces.
@@
If $\beta_{1}=\alpha_{1}+1$, then the above Hölder estimate gives a local Lipschitz estimate, which will imply the existence and some estimate of the gradient for the heat kernel. We give some examples as follows.
  • Unbounded cable systems satisfying V($\Phi$) and HK($\Psi$). We have $\alpha_{1}=1$ and $\beta_{1}=2$ and this gives indeed also Corollary 2.8.
  • The unbounded Vicsek set. We have $\alpha_{1}=\alpha_{2}=\log 5/\log 3$, $\beta_{1}=\beta_{2}=\log 15/\log 3$ and the above Hölder estimate is indeed a Lipschitz estimate, which was also given by Baudoin, Chen [15, Theorem 3.3]. However, a resolvent kernel estimate [5, Theorem 3.40] was intrinsically needed in their proof.
  • The blowups of the Vicsek set by another fractal. We have $\alpha_{1}=\log 5/\log 3$ and $\beta_{1}=\log 15/\log 3$. $\alpha_{2}$ and $\beta_{2}$ will be the Hausdorff dimension and walk dimension of another fractal, say the Sierpiński carpet or the unit interval, see Figure 7 or Figure 8, respectively. The estimates of the gradient for heat kernel can behave differently at small scale and large scale. We will give a detailed discussion in Section 9 about these two examples.
@@ 3 Heat kernel estimates
$\lambda_{1}(D)=\inf\left\{\frac{\mathcal{E}(u,u)}{\lVert u\rVert_{2}^{2}}:u\in\mathcal{F}_{D}\backslash\left\{0\right\}\right\}.$
@@
$\int_{B}|u-u_{B}|^{2}\mathrm{d}m\leq C_{P}\Psi(r)\int_{2B}\mathrm{d}\Gamma(u,u),$ PI($\Psi$)
where $u_{A}$ is the mean value of $u$ on a measurable set $A$ with $m(A)\in(0,+\infty)$, that is,
$u_{A}=\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{A}u\mathrm{d}m=\frac{1}{m(A)}\int_{A}u\mathrm{d}m.$
@@
$\int_{B(x,R+r)\backslash\overline{B(x,R)}}f^{2}\mathrm{d}\Gamma(\varphi,\varphi)$
$\leq\frac{1}{8}\int_{B(x,R+r)\backslash\overline{B(x,R)}}\varphi^{2}\mathrm{d}\Gamma(f,f)+\frac{C_{S}}{\Psi(r)}\int_{B(x,R+r)\backslash\overline{B(x,R)}}f^{2}\mathrm{d}m.$ CS($\Psi$)
We have the characterisations of heat kernel estimates as follows.
Proposition 3.1.
([2, Theorem 1.12]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD. Then the followings are equivalent.
@@
Proposition 3.2.
([31, THEOREM 1.2]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$). Then the followings are equivalent.
@@ 4 Poisson equation on unbounded MMD spaces
First, we have the existence, the uniqueness and the regularity for the solutions of Poisson equation as follows.
Lemma 4.1.
([25, Lemma 2.2, Lemma 2.6]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying FK($\Psi,\nu$). Let $q=\frac{2}{1-\nu}$. Then for any $p\in\left[\frac{q}{q-1},+\infty\right)$, for any ball $B=B(x_{0},r)$, for any $f\in L^{p}(B)$, there exists a unique 222in the sense that if $u_{1},u_{2}\in\mathcal{F}_{B}$ satisfy $\Delta u_{1}=\Delta u_{2}=f$ in $B$, then $u_{1}=u_{2}$ $m$-a.e.. $u\in\mathcal{F}_{B}$ such that $\Delta u=f$ in $B$. There exists $C\in(0,+\infty)$ depending only on $\nu$, $C_{F}$ such that
[…]
Second, we have the pointwise estimate for the solutions of Poisson equation as follows.
Lemma 4.2.
([25, Lemma 2.7]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD, FK($\Psi,\nu$) and CS($\Psi$). Let $q=\frac{2}{1-\nu}$. Then for any $p\in\left[\frac{q}{q-1},+\infty\right)$, there exists $C\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $f\in L^{\infty}(2B)$, if $u\in\mathcal{F}$ satisfies $\Delta u=f$ in $2B$, then for $m$-a.e. $x\in B$, we have
$|u(x)|\leq C\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+F_{1}(x)\right),$
@@
$F_{1}(x)=\sum_{j\leq[\log_{2}r]}\Psi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
The proof of the above two results in [25] was given for the case $\beta_{1}=2$. It is easy to see that the same proof holds for more general $\Psi$, including the case in our paper.
Third, we have Hölder estimate for the solutions of Poisson equation as follows. The proof is similar to [25, Proposition 5.1] and [22, Theorem 3.2] to do gradient estimate. Some technical modifications will be needed.
@@
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u(x)-u(y)|\leq C\left(\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+F_{2}(x)+F_{2}(y)\right),$
@@
$F_{2}(x)=\sum_{j\leq[\log_{2}r]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
@@
[…]
@@
$\lesssim\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|u-u_{k_{1}}|\mathrm{d}m$
$\leq\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|u|\mathrm{d}m+\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|u_{k_{1}}|\mathrm{d}m\right)$
$\lesssim\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+\Psi\left(2^{k_{1}}\right)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+\Phi(2^{k_{1}})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|f|^{p}\mathrm{d}m\right)^{1/p}.$

### v3

@@ 2 Statement of the main results
Let us compare the stability of the gradient-type and Hölder-type conditions as follows.
By a standard argument, under $(UE)$, any of the equivalent gradient-type conditions implies the near-diagonal Gaussian lower bound $(NLE)$ for the heat kernel. In particular, assuming $(UE)$, we have ($\text{GLY}_{\infty}$) implies $(NLE)$. However, the converse is usually not true. Indeed, the conjunction of $(UE)$ and $(NLE)$, or equivalently, HK(2), is equivalent to the conjunction of the volume doubling condition and the scale-invariant $L^{2}$-Poincaré inequality on balls, and is therefore quasi-isometry stable. By contrast, to obtain ($\text{GLY}_{\infty}$), some curvature assumptions are usually needed, see [49, 17], and hence ($\text{GLY}_{\infty}$) is quasi-isometry unstable.
For comparison, by our main result (Theorem 2.1), under UHK($\Psi$), all the Hölder-type conditions are equivalent to NLE($\Psi$). By the quasi-isometry stability of UHK($\Psi$) (see [3]), as well as the quasi-isometry stability of the conjunction of UHK($\Psi$) and NLE($\Psi$) (see [33]), our Hölder-type conditions are quasi-isometry stable. Roughly speaking, this stability stems from the strongly recurrent condition, which forces our MMD spaces to behave analogously to $\mathbb{R}^{1}$ in the classical Riemannian setting.
@@
$(X,d,m,\mathcal{E},\mathcal{F})$ be the corresponding unbounded cable system, equipped with the elementary gradient operator $\nabla$, see [26, Section 3] for more details. Let $\alpha_{1}=1$, $\beta_{1}=2$, $\alpha_{2}=\alpha$, $\beta_{2}=\beta$.
@@
$\lVert{\lvert\nabla u\rvert}\rVert_{L^{\infty}(B)}\leq C_{H}\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m.$ GRH($\Phi,\Psi$)
It is obvious that, on an unbounded cable system, HR($\Phi,\Psi$) implies GRH($\Phi,\Psi$). Indeed, assuming HR($\Phi,\Psi$), the LHS of HR($\Phi,\Psi$) reduces to |u(x)u(y)|/d(x,y)|u(x)-u(y)|/d(x,y) for $d(x,y)\in(0,1)$. Letting $y\to x$ and using the Lebesgue density theorem, we obtain the limit $|\nabla u(x)|$. Taking the essential supremum over $x\in B$, we have GRH($\Phi,\Psi$). Roughly speaking, HR($\Phi,\Psi$) encodes more Hölder continuity information than GRH($\Phi,\Psi$) about the behavior of harmonic functions at large scales.
We give some results as important applications of our main result, Theorem 2.1, to unbounded cable systems. First, since two-sided heat kernel estimates have already been established for a large class of fractals (see [6, 36, 44, 45]), including the Sierpiński carpet (see [9, 10, 37]), we have GRH($\Phi,\Psi$) holds on a large class of fractal-like cable systems, including the Sierpiński carpet cable system, a case whose validity was left open in [26].
@@
HK($\Psi$) with $\alpha_{1}=1$, $\beta_{1}=2$, and $\alpha_{2}=\alpha<\beta=\beta_{2}$. Then GRH($\Phi,\Psi$) holds. In particular, GRH($\Phi,\Psi$) holds on the Sierpiński carpet cable system with $\alpha=\log 8/\log 3<2<\beta\approx 2.09697$.
@@
In [26], the authors verified GRH($\Phi,\Psi$) directly on the Vicsek cable system and the Sierpiński gasket cable system, using certain Hölder regularity [^1] The Hölder regularity follows indeed from a certain extension algorithm for harmonic functions. of harmonic functions on the corresponding fractals [56], which is also available on some class of p.c.f. self-similar sets, but not available on general fractals, especially on the Sierpiński carpet.
Second, we improve the results about gradient estimates for the heat kernel as follows.
@@
HK($\Psi$) with $\alpha_{1}=1$, $\beta_{1}=2$, and $\alpha_{2}=\alpha<\beta=\beta_{2}$.
@@
(1) constitutes an improvement of [26, Theorem 1.1], while (2) improves upon [26, Corollary 1.3], where GRH($\Phi,\Psi$) was assumed. In the present work, we obtain the gradient estimate GHK($\Phi,\Psi$) without imposing gradient-type conditions such as GRH($\Phi,\Psi$).
As a direct consequence of $\text{HHK}_{\text{exp}}(\Phi,\Psi)$, we obtain the following Hölder estimate for the heat kernel on unbounded MMD spaces.
@@
If $\beta_{1}=\alpha_{1}+1$, then the above Hölder estimate gives a local Lipschitz estimate, which will imply the existence and some estimate of the gradient of the heat kernel. We give some examples as follows.
  • Unbounded cable systems satisfying V($\Phi$) and HK($\Psi$). In this case, $\alpha_{1}=1$ and $\beta_{1}=2$. This indeed recovers Corollary 2.8.
  • The unbounded Vicsek set. Here $\alpha_{1}=\alpha_{2}=\log 5/\log 3$ and $\beta_{1}=\beta_{2}=\log 15/\log 3$, hence the above Hölder estimate is indeed a Lipschitz estimate, which was also given by Baudoin, Chen [16, Theorem 3.3]. However, a resolvent kernel estimate [6, Theorem 3.40] was intrinsically needed in their proof.
  • Blowups of the Vicsek set by another fractal. In this case, $\alpha_{1}=\log 5/\log 3$ and $\beta_{1}=\log 15/\log 3$, while $\alpha_{2}$ and $\beta_{2}$ will be given by the Hausdorff dimension and the walk dimension of another fractal, respectively (for instance, the Sierpiński carpet or the unit interval, see Figure 7 and Figure 8). The resulting gradient estimates for the heat kernel can behave differently at small and large scales. We will give a detailed discussion of these two examples in Section 9.
In this paper, we assume a-priori the strongly recurrent condition (1.3), which singles out specific choices of $\Phi,\Psi$. This framework contains many classical fractals, including the Sierpiński gasket, the Sierpiński carpet, nested fractals, and p.c.f. self-similar sets. Starting from this setting, further examples can be constructed via tensor products. For instance, let $X$ be a space with Hausdorff dimension $\alpha$ and walk dimension $\beta>\alpha$. By taking the $n$-fold tensor product $X^{\otimes n}$, one obtains a space with Hausdorff dimension $n\alpha$ and walk dimension $\beta$. For $n$ sufficiently large, we have $n\alpha\geq\beta$, while the Hölder exponent of harmonic functions on $X^{\otimes n}$ remains equal to $\beta-\alpha$, the same as that on $X$; see [2, Proposition 3.8]. For the general case in which that each factor space is equipped with scaling functions $\Phi,\Psi$, see [1, Proposition 3.7].
@@ 3 Heat kernel estimates
[…]
@@
$\int_{B}|u-u_{B}|^{2}\mathrm{d}m\leq C_{P}\Psi(r)\int_{2B}\mathrm{d}\Gamma(u,u).$ PI($\Psi$)
@@
[…]
We have the characterizations of heat kernel estimates as follows.
Proposition 3.1 ([3, Theorem 1.12]).
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD. Then the followings are equivalent.
@@
Proposition 3.2 ([33, THEOREM 1.2]).
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$). Then the followings are equivalent.
@@ 4 Poisson equation on unbounded MMD spaces
First, we have existence, uniqueness, and regularity for solutions of the Poisson equation as follows.
Lemma 4.1 ([26, Lemma 2.2, Lemma 2.6]).
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying FK($\Psi,\nu$). Let $q=\frac{2}{1-\nu}$. Then for any $p\in\left[\frac{q}{q-1},+\infty\right)$, for any ball $B=B(x_{0},r)$, for any $f\in L^{p}(B)$, there exists a unique222in the sense that if $u_{1},u_{2}\in\mathcal{F}_{B}$ satisfy $\Delta u_{1}=\Delta u_{2}=f$ in $B$, then $u_{1}=u_{2}$ $m$-a.e.. $u\in\mathcal{F}_{B}$ such that $\Delta u=f$ in $B$. There exists $C\in(0,+\infty)$ depending only on $\nu$, $C_{F}$ such that
$\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{B}|u|\mathrm{d}m\leq C\sqrt{\Psi(r)}\left(\frac{1}{m(B)}\mathcal{E}(u,u)\right)^{1/2}\leq C\Psi(r)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{B}|f|^{p}\mathrm{d}m\right)^{1/p}.$
Second, we have pointwise estimates for solutions of the Poisson equation as follows.
Lemma 4.2 ([26, Lemma 2.7]).
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD, FK($\Psi,\nu$) and CS($\Psi$). Let $q=\frac{2}{1-\nu}$. Then for any $p\in\left[\frac{q}{q-1},+\infty\right)$, there exists $C\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $f\in L^{\infty}(2B)$, if $u\in\mathcal{F}$ satisfies $\Delta u=f$ in $2B$, then for $m$-a.e. $x\in B$, we have
$|u(x)|\leq C\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{2B}|u|\mathrm{d}m+F_{1}(x)\right),$
@@
$F_{1}(x)=\sum_{j\leq[\log_{2}r]}\Psi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
The proofs of the above two results in [26] were given for the case $\beta_{1}=2$. It is easy to see that the same proofs remain valid for more general $\Psi$, including the case in our paper.
Third, we have Hölder estimates for solutions of the Poisson equation as follows. The proof is similar to that of [26, Proposition 5.1] and [23, Theorem 3.2] for gradient estimates, with some technical modifications.
@@
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u(x)-u(y)|\leq C\left(\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{2B}|u|\mathrm{d}m+F_{2}(x)+F_{2}(y)\right),$
@@
$F_{2}(x)=\sum_{j\leq[\log_{2}r]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
@@
$\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k-1})}|u_{k}|\mathrm{d}m\lesssim\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|u_{k}|\mathrm{d}m\lesssim\Psi\left(2^{k}\right)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
@@
$\lesssim\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{1}})}|u-u_{k_{1}}|\mathrm{d}m$
$\leq\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{1}})}|u|\mathrm{d}m+\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{1}})}|u_{k_{1}}|\mathrm{d}m\right)$
$\lesssim\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m+\Psi\left(2^{k_{1}}\right)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{1}})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m+\Phi(2^{k_{1}})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{1}})}|f|^{p}\mathrm{d}m\right)^{1/p}.$